OpenStudy (anonymous):

The formula for converting degrees Fahrenheit (F) to degrees Celsius (C) is C = (F – 32). Find C when F = 41°

OpenStudy (ghazi):

c=41-32

OpenStudy (amistre64):

forgot the 9/5

OpenStudy (amistre64):

or is it 5/9? i can never remember which one goes to the C

OpenStudy (amistre64):

100 = s(212-32) 100 = s(180) s = 10/18 = 5/9

OpenStudy (amistre64):

.. you forgot the 5/9 :)